# Almost device-independent certification of GME states with minimal measurements

### 28 Feb 2024 v1

@@ Abstract
Device-independent certification of quantum states allows the characterization of quantum states present inside a device by making minimal physical assumptions. A major problem in this regard is to certify quantum states using minimal resources. In this work, we consider the multipartite quantum steering scenario with an arbitrary number of parties but only one of which is trusted in the sense that the measurements performed by the trusted party are known. Consequently, the self-testing scheme is almost device-independent. Importantly, all the parties can only perform two measurements each which is the minimal number of measurements required to observe any form of quantum nonlocality. Then, we propose steering inequalities that are maximally violated by three major classes of genuinely multipartite entangled (GME) states, one, graph states of arbitrary local dimension, two, Schmidt states of arbitrary local dimension, and, three, $N$-qubit generalized W states. Using the proposed inequalities, we then provide an almost device-independent certification of the above GME states.
@@ I Introduction
The strongest DI certification scheme is referred to as self-testing [7] (see also Ref. [8]). It allows one to infer the form of the underlying quantum system solely from the observed measurement outcomes, without needing to know the details of how the measurements were performed. Recently, a lot of attention has been devoted to finding schemes to self-test bipartite quantum states [9, 10, 11, 12, 13] and multipartite quantum states [14, 15, 16, 17, 18, 19, 20, 21, 22]. In particular [20] provides a scheme to self-test any pure entangled quantum state and [21] provides a scheme to self-test any quantum state.
@@
A major problem in this regard is to find the optimal self-testing scheme, that is, using the minimal number of measurements to certify a particular quantum state. As the self-testing schemes are based on quantum nonlocality, each party in the experiment needs to perform at least two measurements. In the bipartite regime, there are only a few schemes that utilize two measurements per party [9, 10, 12] and even fewer in the multipartite scenario [14, 18]. Moreover, in the multipartite scenario most of the schemes are devoted to quantum states that are locally qubits. The main difficulty here consists in designing a suitable Bell inequality that is maximally violated by a given multipartite entangled quantum state. The problem becomes even more complex if one wants to find a Bell inequality which involves the minimal number of two measurements per observer.
@@
To lower the complexity, one can introduce physically motivated assumptions into the considered scenario, leading to what is known as semi-DI certification schemes. One can for instance assume that one of the parties are trusted and their measuring devices perform known measurements. Such an assumption gives rise to scenarios which are typically referred to as one-sided-device-independent (1SDI), and the key resource enabling certification in this case is a weaker form of quantum nonlocality known as quantum steering [23, 24, 25]. Recent investigations have explored the potential of such scenarios for self-testing quantum states and measurements [26, 27, 28, 29, 30, 31, 32, 33]. For this work, we utilise the framework of multipartite quantum steering introduced in [34].
@@
We consider here a much more complicated problem of certifying genuinely multipartite entangled (GME) states shared among an arbitrary number of parties and arbitrary local dimension using minimal number of measurements, that is, two. To simplify the problem we make the aforementioned assumption that only one party is trusted, thus, making our scheme almost device-independent. Furthermore, we restrict the parties to only choose two measurements each. We then construct steering inequalities that are maximally violated by three major classes of GME states. We then prove that maximal violation of our inequalities can be used for certification of the corresponding states. Since, among the arbitrary number of parties only one is trusted, the presented results below are almost DI certificates of the quantum state. The first family of states we can certify in this way is any connected graph state of arbitrary local dimension. Graph states are highly useful in quantum information tasks [35, 36, 37, 38]. The only other known schemes to self-test graph states are restricted to local dimension two [15, 14] and three [19]. The second class of states we consider here are the Schmidt states of arbitrary local dimension, for which a self-testing method was provided in Ref. [17]. The latter, however, requires more measurements per observer and thus is not optimal as far as physical implementations are concerned. Interestingly, our scheme can also be utilized to certify two measurements with the untrusted parties to be mutually unbiased using almost no entanglement across multiple parties. Finally, we self-test the generalized $W$ state (see below for a definition) shared among an arbitrary number of parties but only with local dimension two. While self-testing methods for the standard $N$-qubit $W$ state were already introduced in Refs. [16, 17], here, by making a mild assumption that only one party is trusted, we can significantly generalize these results to a multi-parameter class of genuinely entangled states.
Before proceeding to the results, let us introduce the relevant notions required for this work.
@@ II Preliminaries
We consider a one-sided DI (1SDI) scenario consisting of $N$-parties, namely Alice and $N-1$ Bobs, whom we refer to as $A,B_{1},\ldots,B_{N-1}$. All the parties are located in space-like separated laboratories and share a quantum state $\rho_{A\mathbf{B}}$, where we denote $\mathbf{B}\equiv B_{1}\ldots B_{N-1}$. In this sense, each one of the $N$ parties has a device to measure their share of $\rho_{A\mathbf{B}}$. Since this is a 1SDI scenario, we assume that Alice’s measurement device is trusted, i.e., her measurements are known. Yet, we do not assume that her device performs full tomography over her share of $\rho_{A\mathbf{B}}$. On the other side, the measurement devices that belong to the $N-1$ Bobs are untrusted and we treat them as “black boxes”. The scenario we just described is also referred to as local steering [39], in which the untrusted parties perform local measurements on their respective states to steer the state owned by the trusted party. In this work, we consider that the trusted party performs the measurements given by the $d-$dimensional generalization of the Pauli $Z$ and $X$ measurement
@@
which for ease of notation will be denoted as $Z$ and $X$ respectively. Fig. 1 contains an illustration of the scenario considered here.
We denote the inputs of all the parties $A,B_{1},\ldots,B_{N-1}$ with $x,y_{1},\ldots,y_{N-1}$ respectively and the corresponding outcomes with $a,b_{1},\ldots,b_{N-1}$. For convenience, we chose the notation $\mathbf{y}\equiv y_{1},\ldots,y_{N-1}$ and $\mathbf{b}\equiv b_{1},\ldots,b_{N-1}$. Thus, the correlations in this scenario are captured by the set of probability distributions $\vec{p}=\{p(a,\mathbf{b}|x,\mathbf{{y}})\}$, where $p(a,\mathbf{b}|x,\mathbf{y})$ tells us the probability of getting outcomes $a$ and $\mathbf{b}$ given that the measurements $x$ and $\mathbf{y}$ have been performed. According to Born’s rule, the latter is expressed as
@@
where $M_{x}=\{M_{a|x}\}_{a}$, $N_{y_{i}}^{(i)}=\{N^{(i)}_{b_{i}|y_{i}}\}_{b_{i}}$ are the measurement operators of Alice and Bobs $B_{i}$, respectively. We recall that the measurement operators $M_{a|x}$ and $N^{(i)}_{b_{i}|y_{i}}$ are positive semi-definite and sum to identity on their corresponding Hilbert spaces.
Figure 1: Depiction of the local steering scenario that we consider. A GME state is shared among $N$ parties: one Alice and $N-1$ Bobs. While Alice performs measurements $Z$ and $X$, all the other Bobs perform unknown measurements.
@@
where $\lambda$ is a hidden variable that follows the probability distribution $p(\lambda)$. The hidden variable $\lambda$ determines the state $\rho^{\lambda}_{A}$ received by Alice and also the responses $p(\mathbf{b}|\mathbf{y},\lambda)$ obtained by $N-1$ Bobs. If the set of probability distributions $\{p(a,\mathbf{b}|x,\mathbf{y})\}$ can not be described by the above model, it leads to violation of a linear steering inequality, [$\mathcal{B}\left[\{p(a,\mathbf{b}|x,\mathbf{y})\}\right]\leqslant\beta_{L}$, where $\mathcal{B}$ is a linear combination of $\{p(a,\mathbf{b}|x,\mathbf{y})\}$ and $\beta_{L}$ is the LHS bound , also called classical bound [41].
In this work we express steering inequalities in terms of generalized expectation values. These are defined in terms of $p(a,\mathbf{b}|x,\mathbf{y})$ as
@@
where $\mathbf{B}_{\mathbf{l}|\mathbf{y}}\equiv(B_{1})_{l_{1}|y_{1}}\ldots(B_{N-1})_{l_{N-1}|y_{N-1}}$, $\omega=\mathrm{e}^{2\pi\mathbbm{i}/d}$, and $k,l_{1},\ldots,l_{N-1}\in\{0,\ldots,d-1\}$. If the probabilities $p(a,\mathbf{b}|x,\mathbf{y})$ originate from a quantum measurement, we can write the expectation values as
@@
where $A^{k}_{x}$ is the $k$-th power of the unitary observable $A_{x}$, satisfying $A_{x}^{d}=\mathbb{1}_{d}$, and $(B_{i})_{l_{i}|y_{i}}$ are the Fourier transforms of the measurement operators $N^{(i)}_{b_{i}|y_{i}}$, i.e.,
@@
In what follows, we will present three classes of steering inequalities that are maximally violated by three classes of GME states, namely the $N$-qudit graph states, the $N$-qudit Schmidt states, and the generalized $N$-qubit $W$-states. Recall that a GME state is one that cannot be written as a tensor product of two other pure states corresponding to two disjoint and nonempty subsets of the parties. Formally, a GME state is a state that cannot be written in the form $|\psi\rangle_{A_{1}\ldots A_{N}}=|\alpha\rangle_{A}\otimes|\beta\rangle_{\overline{A}}$ for any bipartition $A|\overline{A}$ such that $A$ is a subset of the parties $A_{1}\ldots A_{N}$, $\overline{A}$ is the complement of $A$ and both $A$ and $\overline{A}$ are non-empty.
@@
Maximally violating steering inequalities plays a critical role in the presented below self-testing results. When experimental results $\vec{p}$ exhibit a maximal violation, they allow us to infer both the observer’s state and the performed measurements up to local isometries. With that in mind, let us formally define multipartite 1SDI self-testing similar to the one introduced in [30]. We assume here that the measurements are projective based on dilation arguments. However, the state $\rho_{N,A\mathbf{B}}$ shared among all parties is in general mixed, which can be purified to $|\psi_{N}\rangle_{A\mathbf{B}E}$ by adding an extra system $E$ as $\rho_{N}=\mathrm{Tr}_{E}(|\psi_{N}\rangle\!\langle\psi_{N}|)$. Alice’s measurements are projective and fixed as $A_{x}$, while Bobs measurements are arbitrary and represented by $B_{i,y_{i}}$. The goal is to certify that the state shared between Alice and Bob is equivalent to a reference state $|\psi^{\prime}\rangle_{A\mathbf{B}^{\prime}}$, as well as, that the measurements performed by Bob are equivalent to the reference measurements $B_{i,y_{i}}^{\prime}$ that generates the same correlation $\vec{p}$. We say that $|\psi\rangle_{A\mathbf{B}E}$ and $B_{i,y_{i}}$ are certified by the statistics $\vec{p}$ if there exists a unitary $U_{i}:\mathcal{H}_{B_{i}}\to\mathcal{H}_{B_{i}}$ such that
@@
$\left(\mathbbm{1}_{A}\otimes\bigotimes_{i=1}^{N-1}U_{i}\right)|\psi_{N}\rangle_{A\mathbf{B}E}$ $=$ $|\psi^{\prime}\rangle_{A\mathbf{B}^{\prime}}\otimes|\phi\rangle_{\mathbf{B}^{\prime\prime}E},$ (7a)
$U_{i}\,B_{i,y_{i}}\,U_{i}^{\dagger}$ $=$ $B_{i,y_{i}}^{\prime}\otimes\mathbbm{1}_{B_{i}^{\prime\prime}},\qquad\forall i$ (7b)
@@
where Bobs Hilbert space is given by $\mathcal{H}_{B_{i}}=\mathcal{H}_{B_{i}^{\prime}}\otimes\mathcal{H}_{B_{i}^{\prime\prime}}$, $|\phi\rangle_{\mathbf{B}^{\prime\prime}E}$ is a junk state in $\mathcal{H}_{\mathbf{B}^{\prime\prime}}\otimes\mathcal{H}_{E}$. Let us remark here that the measurements can only be certified on the local supports of the state, thus, without loss of generality, we assume that the local supports are full-rank.
III Results

### 13 Mar 2026 v2

# Almost device-independent certification of multipartite quantum states with minimal measurements

@@ Abstract
Device-independent certification of quantum states enables the characterization of states within a device under minimal physical assumptions. A major problem in this regard is to certify quantum states using minimal resources. Aiming to address this problem, we consider a multipartite quantum steering scenario involving an arbitrary number of parties, of which only one is trusted, meaning that the measurements performed by this party are known. Consequently, the self-testing scheme is almost device-independent. Importantly, all the parties can only perform two measurements each, which is the minimal number of measurements required to observe any form of quantum nonlocality. Then, we propose steering inequalities that are maximally violated by three major classes of genuinely multipartite entangled (GME) states: graph states of arbitrary local dimension, Schmidt states of arbitrary local dimension, and $N$-qubit generalized W states. Using the proposed inequalities, we then provide an almost device-independent certification of the above GME states. Restricting to qubits, we also lift our almost device-independent scheme to device-independent self-testing.
@@ I Introduction
The most complete form of DI certification is referred to as self-testing [7] (see also Ref. [8]).It allows one to infer the structure of the underlying quantum system solely from the observed Bell-type correlations, without requiring any knowledge of the measurement implementation. Recently, a great deal of attention has been devoted to finding schemes to self-test entangled quantum states in the bipartite [9, 10, 11, 12, 13] and multipartite [14, 15, 16, 17, 18, 19, 20, 21, 22] frameworks. In particular, [20] provides a scheme for self-testing any pure multipartite entangled quantum state and [21] provides a scheme to self-test any quantum state. Although both results are general, they are based on quantum networks which significantly impacts their practical applicability.
@@
A major problem in this regard is to find the optimal self-testing scheme in terms of the number of measurements required to certify a particular quantum state from the observed correlations. Clearly, as self-testing is based on quantum nonlocality, the minimal number of measurements per party is two. In the bipartite regime, there are only a few schemes that utilize two measurements per observer [9, 10, 12] and even fewer schemes exist in the multipartite scenario [14, 18]. Moreover, most of the strategies existing in the multipartite scenario are devoted to quantum states that are locally qubits. One of the main challenges lies in designing a suitable Bell inequality that is maximally violated by a given multipartite entangled quantum state. This task becomes even more complex when one aims to construct such an inequality using only two measurements per observer.
@@
To reduce complexity, one can introduce physically motivated assumptions into the considered scenario, leading to the concept of semi-device-independent certification. For example, one may assume that one of the parties is trusted, meaning that their measurement devices are fully characterized. This leads to scenarios typically referred to as one-sided device-independent (1SDI), where the key resource enabling certification is a weaker form of quantum nonlocality known as quantum steering [23, 24, 25]. Recent investigations have explored the potential of such scenarios for self-testing quantum states and measurements [26, 27, 28, 29, 30, 31, 32, 33]. For this work, we used the multipartite quantum steering framework introduced in [34].
@@
Here, we address a significantly more challenging problem: certifying genuinely multipartite entangled (GME) states shared among an arbitrary number of parties with arbitrary local dimensions, using only a minimal number of measurements, namely, two per party. To simplify the problem we make the aforementioned assumption that a single party is trusted, thus, making our scheme almost device-independent. Furthermore, we restrict the parties to only choose two measurements each. Under these assumptions, we construct steering inequalities that are maximally violated by three relevant classes of GME states. We then prove that maximal violation of our inequalities can be used for certification of the corresponding states. Since, among the arbitrary number of parties only one is trusted, the presented results below are almost DI certificates of the quantum state. The first family of states we can certify in this way are graph states of arbitrary local dimension. Graph states form one of the most representative classes of states, which includes for instance the well-known Greenberger-Horne-Zeilinger (GHZ) [35], cluster [36] or the absolutely maximally entangled (AME) [37] states. Graph states are key resources for many applications, just to mention measurement-based quantum computing [38, 39], quantum metrology [40] or multipartite secret sharing [41, 42]. The existing self-testing schemes for graph states are restricted to local dimension two [15, 14] and three [19]. Here we go beyond these particular dimensions and provide strategies that allow one to certify graph states of arbitrary local dimension, yet at the cost that one of the measurement devices is trusted.
The second class of states we consider here are the Schmidt states of arbitrary local dimension, which also include the aforementioned GHZ states. The Schmidt states are useful for quantum metrology [43] or multipartite secret sharing [41]. A self-testing method for this class was provided in Ref. [17]. The latter, however, requires more measurements per observer and thus is not optimal as far as physical implementations are concerned. Interestingly, our scheme can also be utilized to certify two measurements with the untrusted parties to be mutually unbiased using almost no entanglement across multiple parties.
We finally provide a self-testing scheme for the generalized $W$ states (see below for a definition) shared among an arbitrary number of parties but only with local dimension two. The generalized $W$ states belong together with the Schmidt states to a broader class of entangled symmetric multipartite states and find applications for instance in quantum metrology [42]. While self-testing methods for the standard $N$-qubit $W$ state were already introduced in Refs. [16, 17], here, by making a mild assumption that only one party is trusted, we can significantly generalise these results to a multi-parameter class of genuinely entangled states.
Finally, we show how the above results can be extended to the device-independent regime by taking inspiration from Ref. [20]. More precisely, we consider an extra party which enables certification of the measurements of the trusted party from violation of the CHSH Bell inequality.
@@ II Preliminaries
We consider a one-sided DI (1SDI) scenario consisting of $N$-parties, AA and $B_{1},\ldots,B_{N-1}$, which are also called Alice and Bobs. All the parties are located in space-like separated laboratories and share a quantum state $\rho_{A\mathbf{B}}$, where we denote $\mathbf{B}\equiv B_{1}\ldots B_{N-1}$. Each party has a measurement device performing measurements on their shares of $\rho_{A\mathbf{B}}$. We additionally assume that Alice’s measurement device is trusted, meaning that it performs known measurements. Yet, we do not assume that her device performs full tomography over her share of $\rho_{A\mathbf{B}}$. On the other hand, the measurement devices that belong to N1N-1 Bobs are untrusted and we treat them as “black boxes”. The scenario we just described is also called local steering [44], in which the untrusted parties perform local measurements in their respective states to steer the state owned by the trusted party.
In this work, we consider that the trusted party performs the measurements given by the $d$-dimensional generalization of the Pauli $Z$ and $X$ observables,
@@
which, for ease of notation, will be denoted as $Z$ and $X$, respectively. Fig. 1 illustrates the scenario considered.
We denote the inputs and outputs of the parties $A$ and $B_{1},\ldots,B_{N-1}$ with $x$ and $y_{1},\ldots,y_{N-1}$ and $a$ and $b_{1},\ldots,b_{N-1}$, respectively. For convenience, we use the notation $\mathbf{y}\equiv y_{1},\ldots,y_{N-1}$ and $\mathbf{b}\equiv b_{1},\ldots,b_{N-1}$. The correlations observed by the parties when they repeat their measurements many times are captured by the set of probability distributions
$\vec{p}=\{p(a,\mathbf{b}|x,\mathbf{{y}})\},$ (2)
where $p(a,\mathbf{b}|x,\mathbf{y})$ stands for the probability of obtaining results $a$ and $\mathbf{b}$ given that the measurements $x$ and $\mathbf{y}$ have been carried out. According to Born’s rule, the latter is expressed as
@@
where $M_{a|x}$ and $N^{(i)}_{b_{i}|y_{i}}$ are the measurement operators representing the measurements of Alice and all Bobs $B_{i}$, respectively; recall that they are positive semi-definite and sum to identity on their corresponding Hilbert spaces.
Figure 1: Depiction of the local steering scenario that we consider. A GME state is shared among NN parties: one Alice and N1N-1 Bobs. While Alice performs measurements $Z$ and $X$, all the other Bobs perform unknown measurements.
@@
where $\lambda$ is a hidden variable that follows the probability distribution $p(\lambda)$. The hidden variable $\lambda$ determines the state $\rho^{\lambda}_{A}$ received by Alice and also the responses $p(\mathbf{b}|\mathbf{y},\lambda)$ obtained by N1N-1 Bobs. If the set of probability distributions $\{p(a,\mathbf{b}|x,\mathbf{y})\}$ cannot be described by the above model, it leads to violation of the linear steering inequality $\mathcal{B}\left[\{p(a,\mathbf{b}|x,\mathbf{y})\}\right]\leqslant\beta_{L}$, where $\mathcal{B}$ is a linear combination of $\{p(a,\mathbf{b}|x,\mathbf{y})\}$ and $\beta_{L}$ is the LHS bound, also called the classical bound [46], which is the maximal value of the functional $\mathcal{B}$ over the LHS models (4).
In this work, we express steering inequalities in terms of generalized expectation values, which are defined through $p(a,\mathbf{b}|x,\mathbf{y})$ in the following way
@@
where $\mathbf{B}_{\mathbf{l}|\mathbf{y}}\equiv(B_{1})_{l_{1}|y_{1}}\ldots(B_{N-1})_{l_{N-1}|y_{N-1}}$ with $k,l_{1},\ldots,l_{N-1}\in\{0,\ldots,d-1\}$, and $\omega=\exp(2\pi\mathbbm{i}/d)$. If the probabilities $p(a,\mathbf{b}|x,\mathbf{y})$ express through Eq. (3), we can write the expectation values as
@@
where $A^{k}_{x}$ is the $k$-th power of the unitary observable $A_{x}$, which, to recall, we assume to be the generalized Pauli observables $Z$ and $X$. Then, $(B_{i})_{l_{i}|y_{i}}$ are operators representing the measurements of $B_{i}$, which are defined through the Fourier transforms of the measurement operators $N^{(i)}_{b_{i}|y_{i}}$, i.e.,
@@
In what follows, we will present new steering inequalities that are maximally violated by three classes of GME states: the $N$-qudit graph states, the $N$-qudit Schmidt states and the generalized $N$-qubit $W$ states. Recall that a GME state is one that cannot be written as a tensor product of two other pure states corresponding to an arbitrary bipartition of all the parties into two disjoint and non-empty subsets. More precisely, an $N$-partite state $|\psi_{A\mathbf{B}}\rangle$ is GME if $|\psi_{A\mathbf{B}}\rangle\neq|\alpha\rangle_{Q}\otimes|\beta\rangle_{\overline{Q}}$ for any two states $|\alpha\rangle_{Q}$, $|\beta\rangle_{\overline{Q}}$ and any bipartition of the parties $A$ and $B_{1},\ldots,B_{N-1}$ into two non-empty and disjoint sets $Q$ and $\overline{Q}$.
@@
Maximal violations of steering inequalities play a critical role in the self-testing results presented below. When the observed correlations $\vec{p}$ exhibit a maximal violation, they allow us to infer both the observer’s state and the performed measurements up to local isometries. With that in mind, let us formally define multipartite 1SDI self-testing similar to the one introduced in [30]. We assume here that the measurements of the untrusted parties are projective based on dilation arguments. At the same time, we do not assume the shared state $\rho_{A\mathbf{B}}$ to be purifiable by the unstrusted Bobs as due to the Schmidt decomposition between $A$ and $\mathbf{B}$ this would imply that the dimension of $\mathbf{B}$’s joint Hilbert space is the same as that of Alice, i.e., $d$. Yet, we can always consider a purification $|\psi\rangle_{A\mathbf{B}E}$ with an extra system $E$ such that $\rho_{A\mathbf{B}}=\mathrm{Tr}_{E}(|\psi\rangle\!\langle\psi|_{A\mathbf{B}E})$, which is not possessed by any of the parties. Alice’s measurements are projective and fixed as $A_{x}$, while the measurements of Bobs are arbitrary and represented by the operators $(B_{i})_{l_{i}|y_{i}}$. The goal is to certify that the state shared between Alice and Bob and the measurements performed by Bobs are equivalent, up to certain equivalences, to a reference state $|\psi^{\prime}\rangle_{A\mathbf{B}^{\prime}}\in(\mathbbm{C}^{d})^{\otimes N}$ and to the reference measurements represented by $(B_{i})_{l_{i}|y_{i}}^{\prime}$ defined on $\mathbbm{C}^{d}$, respectively, which all give rise to the same correlations $\vec{p}$. More formally, we say that $|\psi^{\prime}\rangle_{A\mathbf{B}^{\prime}}$ and $(B^{\prime}_{i})_{l_{i}|y_{i}}$ are certified from the observed $\vec{p}$ if there exists a unitary $U_{i}:\mathcal{H}_{B_{i}}\to\mathcal{H}_{B_{i}}$ for each Bob $B_{i}$ such that
@@
$\left(\mathbbm{1}_{A}\otimes\bigotimes_{i=1}^{N-1}U_{i}\right)|\psi\rangle_{A\mathbf{B}E}=|\psi^{\prime}\rangle_{A\mathbf{B}^{\prime}}\otimes|\phi\rangle_{\mathbf{B}^{\prime\prime}E},$ (8)
and
$U_{i}\,(B_{i})_{l_{i}|y_{i}}\,U_{i}^{\dagger}=(B^{\prime}_{i})_{l_{i}|y_{i}}\otimes\mathbbm{1}_{B_{i}^{\prime\prime}},\qquad\forall i$ (9)
@@
where Bob’s Hilbert spaces decompose into tensor products $\mathcal{H}_{B_{i}}=(\mathbbm{C}^{d})_{B_{i}^{\prime}}\otimes\mathcal{H}_{B_{i}^{\prime\prime}}$, where $\mathcal{H}_{B_{i}^{\prime\prime}}$ is an auxiliary Hilbert space of unknown dimension and $|\phi\rangle_{\mathbf{B}^{\prime\prime}E}$ is a junk state from $\mathcal{H}_{\mathbf{B}^{\prime\prime}}\otimes\mathcal{H}_{E}$. Let us remark here that the measurements can only be certified on the local supports of the state, thus, without loss of generality, we assume that the local density matrices of the joint state $|\psi\rangle_{A\mathbf{B}E}$ are full-rank.
III 1SDI certification of multipartite quantum states